# Project Page: https://yongtaoge.github.io/projects/humanwild

### v2

@@ 4.5 Comparing with Other Real/Synthetic Datasets
In Table 6, we analyze the performance of different training datasets using mixed strategies on CLIFF with HRNet-W48 backbone. Key findings are: 1). CG-rendered data outperforms real-world data (motion-capture, pseudo-labeled) on 3DPW and RICH datasets, suggesting synthetic data advantages. 2). Solely training on HumanWild yields slightly lower performance than BEDLAM, we conjecture that there still exists label noise, especially in hand and face parameters. 3). Combining HumanWild with CG-rendered datasets (BEDLAM, AGORA) and marker-based dataset (3DPW) enhances performance, emphasizing the benefit of diverse data integration.
Table 6: Reconstruction error on 3DPW and RICH. ‘R’ denotes mixed real-world datasets, ‘H’ denotes $\tt HumanWild$, ‘B’ denotes BEDLAM [5], and ‘A’ denotes AGORA [55]. ‘P’ denotes 3DPW [46].
@@
CLIFF[39, 5] H+B+A+P 41.9 65.2 76.8 48.4 79.7 91.1
@@ 4.6 Result on Challenging Benchmarks
To verify the effectiveness of $\tt HumanWild$ on different scenerios, we report the detailed performance on three challenging benchmarks. Firstly, we evaluate the capability of $\tt HumanWild$ in handling perspective distortion tasks, by performing fair ablation using Zolly [77] with ResNet50 [16] backbone. The results in Tab. 8 show that $\tt HumanWild$ dataset achieves higher generalization performance on SPEC-MTP [32] when compared with PDhuman [77], which is a CG-rendered perspective distortion dataset. Secondly, MOYO [71] encompasses challenging yoga poses not present in BEDLAM and $\tt HumanWild$ datasets, providing a unique testbed for evaluating the model’s generalization ability on hard poses. The results in Tab. 8 demonstrate that $\tt HumanWild$, complements existing datasets, resulting in improved generalization performance on the MOYO test set. In Tab. 9, our analysis reveals that while Humanwild demonstrates superior performance over BEDLAM on body metrics, it exhibits relatively poorer performance on hand and face metrics, particularly evident in comparison to the AGORA test set. We hypothesize that the alignment of hand and facial annotations with generated images remains noisy, primarily due to inherent limitations in current diffusion models’ ability to accurately generate hand poses, particularly within small resolutions.
Table 7: SPEC-MTP.
Table 8: MOYO.
Training Data
[…]
PDHuman
102.4
159.1
168.0
$\tt HumanWild$
90.0 151.8 160.9
Training Data
MPJPE $\downarrow$ PA-MPJPE $\downarrow$ PVE $\downarrow$ PA-PVE $\downarrow$
B+A+P
143.1 94.9 163.9 95.7
B+A+P+H
134.6 89.1 154.1 88.9
Table 8: MOYO.
Table 9: AGORA test set. normal-$\dagger$ denotes the methods that are finetuned on the AGORA training set. normal-$\ast$ denotes the methods that are trained on AGORA training set only.
@@
Hand4Whole [50$\dagger$ 144.1 96.0 141.1 92.7 135.5 90.2 41.6 46.3 48.1 132.6 87.1 46.1 44.3 46.2
BEDLAM [5$\dagger$ 142.2 102.1 141.0 101.8 103.8 74.5 23.1 31.7 33.2 102.9 74.3 24.7 29.9 31.3
PyMaF-X [87$\dagger$ 141.2 94.4 140.0 93.5 125.7 84.0 35.0 44.6 45.6 124.6 83.2 37.9 42.5 43.7
OSX [40$\ast$ 130.6 85.3 127.6 83.3 122.8 80.2 36.2 45.4 46.1 119.9 78.3 37.9 43.0 43.9
$\tt HumanWild$dagger 120.5 73.7 115.7 72.3 112.1 68.5 37.0 46.7 47.0 107.6 67.2 38.5 41.2 41.4
@@ 5 Discussion and Conclusions
Based on the experiments, we can answer the question “Is synthetic data generated by generative models complementary to CG-rendered data for the 3D HPS task?” Our results strongly suggest that through the integration of 3D human priors with ControlNet, it becomes feasible to produce high-fidelity pseudo labels encompassing a wide array of real-world scenarios. As the landscape sees the emergence of increasingly large generative models, there arises a promising prospect for the expansion of diverse 3D human training datasets without intricate mocap systems and CG pipelines. We hope that our endeavors could pave the way for leveraging generative models to generate high-quality datasets conducive to enhancing the efficacy of 3D human perception tasks.
@@
Our pipeline can apply to a series of similar tasks where high-quality data pairs are hard to collect, e.g., 3D animal pose estimation and 3D reconstruction of human-object/human-human interaction. Addressing these challenges requires further advancements in our methodology.
Acknowledgement
This work was supported by National Key R&D Program of China (No. 2022ZD0118700).
Appendix 0.A Appendix
0.A.1 Implement Details for training ControlNet
@@
$\tt HumanWild$ employs a customized ControlNet to generate human images and corresponding annotations. In the preliminary phase of our research, we adopt the off-the-shelf depth-to-image ControlNet, and keypoint-to-image ControlNet, which are trained on the LAION5B dataset without filtering low-quality images. Experiments show that the pre-trained models face challenges when it comes to generating human images with hard poses. To resolve this issue, we collect high-quality human images from multiple datasets and then conduct finetuning on the collected datasets for better generation performance.
Data preprocessing.
We curate from four datasets LAION5B [64], COYO [7], HI4D [86] and BEDLAM [5]. Following HyperHuman [43], we employ YOLOS [13] for human detection on LAION5B [64] and COYO-700M [7] datasets, images containing 1 to 3 human bounding boxes are retained. We use BLIP [37] to get text captions for BEDLAM [5] and HI4D [86].
Training.
We train the customized ControlNet with with AdamW [29] in 1e55-5- 5 learning rate, and 0.01 weight decay for $80,000$ iterations on 8 A100 GPUs.
0.A.2 Failure Cases of the keypoint and depth-based ControlNet
@@
In Fig. 6, we show some failure cases of the initial data pairs generated by a multi-condition, i.e., depth and keypoint-based, ControlNet The inconsistency of the image and 3D mesh would affect the performance of the 3D human pose estimation. Thus, it is necessary to retrain a customized ControlNet with the surface normal condition proposed in this work.
@@
Figure 6: Failure cases of the initial data pairs generated by ControlNet. See highlighted regions with red circles.
0.A.3 More Visualizations of Human Interactions
We show visualization results on human interactions. Fig. 7 demonstrates that our pipeline can generate well-aligned image-annotation pairs where people are with close interactions. The generated data pairs are of great value in enhancing existing human interaction datasets collected in the studio environment. (e.g., Hi4D [86] and CHI3D [14])
@@ 0.A.3 More Visualizations of Human Interactions
0.A.4 Text prompt examples generated by LLM
In Tab. 10, we show some text prompt examples, which are generated by ChatGPT [52] with diverse human actions and scenes.
Table 10: Text prompt examples.
@@ 0.A.4 Text prompt examples generated by LLM
0.A.5 Misalignment Analysis of Complicated Prompts and Normals
We present instances of failure where text prompts conflict with the surface normal condition. Specifically, when the text prompt suggests an action that deviates from the surface normal map, the resulting images often fail to adhere to the text prompt, particularly when the control factor of the surface normal map approaches $1.0$. As shown in Fig. 8, despite employing the same input surface normal map and a control factor of $0.95$, the generated images exhibit similar foreground human identities but differ in background elements corresponding to distinct text prompts. Fig. 9 illustrates that when the control factor of the surface normal is extremely low, the generated images may disregard the surface normal map condition.
0.A.6 Results on 2D HPE
2D keypoint refinement. $\tt HumanWild$ can also be applied for the 2D human pose estimation task. Specifically, we obtain the initial 2D keypoints by projecting the 3D joints of SMPL-X into the image plane with the simulated camera parameters. The intuition of the 2D keypoint refinement is that different pose datasets provide different skeleton formats [63], even though they sometimes share the same joint names. To tackle the label discrepancies, it is necessary to refine the initial 2D keypoints to the formulation of the target 2D pose dataset. Here, we take the COCO dataset as an example to explain the proposed strategy for refining the initial 2D keypoints converted from the SMPL-X model.
We leverage a COCO pre-trained keypoint decoder proposed in [45] to get more accurate 2D keypoint labels. Concretely, we replace the coarse proposals from fully connected layers with the initial 2D keypoints converted from SMPL-X, and then several deformable cross-attention [91] operations are performed between the image features and keypoint queries to gradually generate the 2D keypoints in the COCO format. Compared to the pure pseudo-labeling process, our refinement strategy has more reliable initial keypoint proposals. Thus, our method has a higher upper bound of the final generated 2D keypoint labels.
In Tab. 12, we adopt two types of 2D pose estimators, RTMPose [23] and RLEPose [35], to verify the effectiveness of the proposed data generation pipeline. For a fair comparison, all the models are trained with 10 epochs. Our pipeline can consistently improve the detection performance when mixed with different COCO training subsets (from $1\%$ to $100\%$). The performance of $\tt HumanWild$ is comparable with BEDLAM in all data crops. When joint training with all three datasets, both 2D pose regressors get the best performance.
We summarize the key results in Tab. 11. (1) Due to the lack of occlusion and multi-person scenes in the generated images, $\tt HumanWild$ cannot improve the results on the OCHuman validation set. (2) $\tt HumanWild$ can outperform DatasetDM [80] by a large margin on the COCO validation set under the same training setting. We conjecture that the generated dataset only has one person per image, which lacks human-scene occlusion and human-human interaction. We also find that classification-based RTMPose is less data-hungry than regression-base RLEPose in low data regime, e.g., achieving much higher AP on $1\%$ COCO training set.
Table 11: Main Results on 2D Human Pose Estimation. ‘P’ denotes $\tt HumanWild$; ‘D’ denotes DatasetDM [80], ‘C’ denotes COCO; ‘B’ denotes BEDLAM [5]. Crops % only applies to COCO during the training. We evaluate results on the COCO and OCHuman Datasets.
Method Backbone Training Set Crop COCO OCHuman
[…]
RTMPose [23] CSPNeXt C 100 75.2 71.6 81.9 69.9 67.0 69.8
RTMPose [23] CSPNeXt P+C 100 75.7 72.4 82.9 67.2 62.5 67.2
SimplePose [81] HRNet-W32 C 100 74.9 71.3 81.5 59.8 65.3 59.8
SimplePose [81] HRNet-W32 D+C 1 47.5 44.2 52.6 N/A N/A N/A
SimplePose [81] HRNet-W32 P+C 1 50.3 44.7 59.1 29.5 18.7 29.5
Figure 10: More Visualization of $\tt HumanWild$ with background surface normal information extracted from ScanNet++ [85]. The first two columns are compositional meshes with both SMPL-X  [56] model and indoor background meshes. The third column is the surface normal extracted from the compositional meshes. The fourth and fifth columns are images generated from the third column. The sixth column is the surface normal map without background normal. The last two columns are images generated from the sixth column.
Table 12: Ablation experiments on 2D human pose estimation. ‘C’ denotes COCO; ‘B’ denotes BEDLAM; ‘P’ denotes $\tt HumanWild$, and Crops % only applies to COCO. All experiments are evaluated on the COCO validation set. AP is used as the evaluation metric.
[…]
RTMPose C Classification CSPNeXt 0.0 7.2 23.6 67.9
RTMPose B+C Classification CSPNeXt 46.4 55.9 58.7 68.4
RTMPose P+C Classification CSPNeXt 49.1 55.7 58.0 68.1
RTMPose P+B+C Classification CSPNeXt 61.9 63.1 64.4 71.3
RLEPose C Regression ResNet50 0.0 3.9 19.2 53.5
RLEPose B+C Regression ResNet50 40.6 47.7 55.3 64.8
RLEPose P+C Regression ResNet50 31.5 39.0 50.3 65.1
RLEPose P+B+C Regression ResNet50 51.8 56.3 58.5 66.6

### v3

@@ 4.6 Comparing with Other Real/Synthetic Datasets
In Table VIII, we analyze the performance of different training datasets using mixed strategies on CLIFF with HRNet-W48 backbone. Key findings are: 1). Pure CG-rendered dataset outperforms real-world data (motion-capture, pseudo-labeled) on 3DPW and RICH datasets, suggesting the effectiveness of the synthetic data. 2). Solely training on HumanWild yields slightly lower performance than BEDLAM; we conjecture that there still exists some levels of label noise, especially in hand and face poses. 3). Combining HumanWild with CG-rendered datasets (BEDLAM, AGORA) and marker-based dataset (3DPW) enhances performance, which demonstrates the flexibility of HumanWild, i.e., can be integrated with existing datasets to boost the estimation performance.
@@
CLIFF[4, 19] H+B+A+P 41.3 65.2 76.7 48.4 79.7 91.1
CLIFF[4, 19] H+B+A+P+R 45.5 71.3 84.3 52.6 86.4 98.3
TABLE VIII: Reconstruction error on 3DPW and RICH. ‘R’ denotes mixed real-world datasets, ‘H’ denotes HumanWild, ‘B’ denotes BEDLAM [19], and ‘A’ denotes AGORA [32]. ‘P’ denotes 3DPW [23].
@@ 4.7 Result on Challenging Benchmarks
To verify the effectiveness of HumanWild on different scenerios, we report the detailed performance on three challenging benchmarks. Firstly, we evaluate the capability of HumanWild in handling perspective distortion human pose and shape estimation, by performing fair ablation using Zolly [56] with ResNet50 [77] backbone. The results in Table˜V show that HumanWild dataset achieves higher generalization performance on SPEC-MTP [76] when compared with PDhuman [56], which is a CG-rendered perspective distortion dataset. Secondly, MOYO [13] encompasses challenging yoga poses not present in BEDLAM and HumanWild datasets, providing a unique testbed for evaluating the model’s generalization ability on hard poses. The results in Table˜V demonstrate that HumanWild, complements existing datasets, resulting in improved generalization performance on the MOYO test set.
We show AGORA test set results in Table˜IX, our analysis reveals that while Humanwild demonstrates superior performance over BEDLAM on body metrics, it exhibits relatively poorer performance on hand and face metrics, particularly evident in comparison to the AGORA test set. We hypothesize that the alignment of hand and facial annotations with generated images remains noisy, primarily due to inherent limitations in current diffusion models’ ability to accurately generate hand poses, particularly within small resolutions.
@@
[…]
HumanWild $\dagger$ 120.5 73.7 115.7 72.3 112.1 68.5 37.0 46.7 47.0 107.6 67.2 38.5 41.2 41.4
TABLE IX: AGORA test set. $\dagger$ denotes the methods that are finetuned on the AGORA training set. $\ast$ denotes the methods that are trained on AGORA training set only.
4.8 Visualization on in-the-wild scenes
In this section, we show visualization results of two types of CLIFF estimators with the same HRNet-W48 backbone [70]. One CLIFF is trained on the BEDLAM dataset (we use the official weights released by authors), annother is trained on both BEDLAM and HumanWild datasets. As we can see from Fig.˜4, our finetuned CLIFF estimator has more accurate pose and shape estimation results in diverse scenes, demonstrating the effectiveness of our data to help existing models have better generalization performance.
5 Limitations and Future Work
Limitations in photo-realism. We observe that our fine-tuned ControlNet model can generate images that are either blurry or lack photorealism. This issue stems from that the underlying SDXL base model possesses limited capacity for photorealistic generation. As shown in Fig.˜8, the SDXL base model tends to generate images with a blurred background. Given recent advancements, such as the FLUX [25] model, which demonstrate more promising results in synthesizing high-resolution and realistic human imagery, future efforts should be directed toward scaling up controllable generative models to achieve greater photorealism.
Figure 8: Images generated directly from SDXL [24]. Text prompt: a beautiful woman wearing a loose dress on the beach.
Limitations in multi-Human synthesis. The current pipeline effectively generates high-quality images containing one to three human subjects. When the number of individuals increases, their scale within the image decreases, leading to poor alignment with the inputs of the guiding depth, keypointmap. Consequently, the resulting data is unsuitable for training human pose and shape models. Future research could leverage recent generative image editing models, such as Bagel [80] and Flux Kontext [81], to translate our computer-rendered datasets into more photorealistic domains.
@@ 6 Discussion and Conclusions
Based on the experiments, we can answer the question “Is synthetic data generated by generative models complementary to CG-rendered data for the 3D HPS task?” Our results strongly suggest that through the integration of 3D human priors with pretrained diffusion models, it becomes feasible to produce high-fidelity pseudo labels encompassing a wide array of real-world scenarios. As the landscape sees the emergence of increasingly large generative models, there arises a promising prospect for the expansion of diverse 3D human training datasets without intricate mocap systems and CG pipelines. We hope that our endeavors could pave the way for leveraging generative models to generate high-quality datasets, which are conducive to enhancing the efficacy of 3D human perception tasks.
@@
Our pipeline holds significant potential for various related tasks where acquiring high-quality data pairs is difficult, such as 3D animal pose estimation and 3D reconstruction of human-object or human-human interactions. Addressing these challenges will require further advancements in our methodology, which we will explore in future research.
[Uncaptioned image] Yongtao Ge Yongtao Ge is currently a PhD student at the University of Adelaide. He has published papers on top-tier conferences, including SIGGRAPH, ICCV, ECCV, ICLR and AAAI. His research interests mainly lies in the intersection of computer vision and computer graphics, including human-centered digitization and world models.
[Uncaptioned image] Wenjia Wang is currently pursuing his PhD degree in The Univeristy of Hongkong, advised by Prof.Taku Komura and Prof. Wenping Wang. Prior to that, he worked in computer vision in Sensetime and Tencent, from 2019 to 2022. He has published papers on top-tier conferences, including CVPR, ECCV, ICCV, and IJCAI.
[Uncaptioned image] Yongfan Chen was a master student at Zhejiang University, under the supervision of Prof. Chunhua Shen. Prior to this, he earned his bachelor’s degree in Computer Science and Technology from Huazhong University of Science and Technology. His current research primarily focuses on computer vision, pose estimation and generative models.
[Uncaptioned image] Fanzhou Wang is currently an Algorithm Researcher at SenseTime Group Inc. Prior to that, she received her M.S. degree and B.E. degree from Southeast University in 2023 and 2020. Her research interests include synthetic data and computer graphics.
[Uncaptioned image] Lei Yang is currently a Research Director at SenseTime Group Inc., under the supervision of Prof. Xiaogang Wang. Prior to that, Lei received his PhD degree from the Chinese University of Hong Kong in 2020, advised by Prof. Dahua Lin and Prof. Xiaoou Tang. Before that, Lei obtained his B.E. degree from Tsinghua University in 2015. He has published over 10 papers on top conferences in relevant fields, including CVPR, ICCV, ECCV, AAAI, SIGGRAPH and RSS.
[Uncaptioned image] Hao Chen is an Assistant Professor of Computer Science at Zhejiang University. Prior to that, he worked as a researcher at Noah’s Ark Lab at Huawei. He got his PhD degree from University of Adelaide. He obtained his undergraduate and master degrees from Zhejiang University.
Chunhua Shen is a Chair Professor of Computer Science at Zhejiang University and Zhejiang University of Technology. He was with Amazon Australia, and The University of Adelaide, Australia. His research interest is Computer Vision and Machine Learning.
Appendix A
A.1 Implement Details for Training ControlNet
@@
HumanWild employs a customized multi-condition ControlNet to generate human images and corresponding annotations. In the preliminary phase of our research, we adopt the off-the-shelf depth-to-image ControlNet, and keypoint-to-image ControlNet, which are trained on the LAION5B dataset without filtering low-quality images. Experiments show that the pre-trained models face challenges when it comes to generating human images with hard poses. To resolve this issue, we collect high-quality human images from multiple datasets and then conduct finetuning on the collected datasets for better generation performance. We train the customized ControlNet with with AdamW [82] in 1e5-5 learning rate, and 0.01 weight decay for $80,000$ iterations on 8 A100 GPUs. For training efficiency, we utilize Fully Sharded Data Parallel (FSDP) with ZeRO Stage 2, gradient checkpointing, and mixed-precision training
Appendix B
B.1 Failure Cases of the Off-the-shelf ControlNet
@@
In Fig.˜11, we show some failure cases of the data pairs generated by off-the-shelf ControlNet trained on general datasets. The inconsistency of the image and 3D mesh would affect the performance of the 3D human pose estimation. Thus, it is necessary to retrain a customized multi-condition ControlNet with the human-centeric dataset collected in this work.
@@
Figure 11: Failure cases of the data pairs generated by general ControlNet. See highlighted regions with red circles.
B.2 More Visualization of Generated Data Sample
We show more data sample visualization with diverse scenes and body shapes in  Fig.˜12.
Figure 12: More Data sample visualization of HumanWild. For each data sample, the left side is the normal image rendered from SMPL-X model, the right side is the image generated by HumanWild pipeline. The first two rows demonstrate the indoor activities. The third row shows the generated images with huge distortion. The fourth row describes the diverse camera views. The fifth row shows overweight human bodies.
@@ B.2 More Visualization of Generated Data Sample
We show visualization results on human interactions in Fig.˜13, which demonstrates that our pipeline can generate well-aligned image-annotation pairs where people are with close interactions. The generated data pairs are of great value in enhancing existing human interaction datasets collected in the studio environment. (e.g., Hi4D [83] and CHI3D [84])
B.3 Text Prompt Examples Generated by LLM
In Table˜X, we show some text prompt examples, which are generated by ChatGPT [85] with diverse human actions and scenes.
TABLE X: Text prompt examples.
@@ B.3 Text Prompt Examples Generated by LLM
B.4 Misalignment Analysis of Complicated Prompts and Spatial Conditions
We present instances of failure where text prompts conflict with the spatial condition, e.g., surface normal. Specifically, when the text prompt suggests an action that deviates from the surface normal map, the resulting images often fail to adhere to the text prompt, particularly when the control factor of the surface normal map approaches $1.0$. As shown in Fig.˜9, despite employing the same input surface normal map and a control factor of $0.95$, the generated images exhibit similar foreground human identities but differ in background elements corresponding to distinct text prompts. Fig.˜10 illustrates that when the control factor of the surface normal is extremely low, the generated images may disregard the surface normal map condition.